OpenStudy (vishweshshrimali5):

The digit in the unit's place of 7^171 + (177)! is (a) 3 (b) 2 (c) 1 (d) 0

OpenStudy (anonymous):

177! is going to have a bunch of zeros at its end, so we just need to figure out what the ones digit of 7^171 is. Notice that: \[7^1 = 7, 7^2 = 49, 7^3 = 343, 7^4 = 2401, 7^5 = 16807, \ldots \] The ones place repeats after the 4th term. What we have is a sequence that goes like: 7, 9, 3, 1, 7, 9, 3, 1,... This sequence repeats every 4 terms, so lets see what the 171th (the power) term is: 171/4 = 42 with a remainder of 3. That means going out 171 terms will cycle through the 4 terms 42 times, then stop on the 3rd term, which is 3. So 7^171 will end with a 3, and likewise: \[7^{171}+(177!)\] will also end in a 3.

OpenStudy (vishweshshrimali5):

Right boss You are a genius